A spinner has 20 equally sized sections, 8 of which are green and 12 of which are red. The spinner is spun and, at the same time, a fair coin is tossed. What is the probability that the spinner lands on red and the coin toss is tails?

Answer by jim_thompson5910(35256)

P(Red AND Tails) = P(Red)*P(Tails)

P(Red AND Tails) = (12/20)*(1/2)

P(Red AND Tails) = (12*1)/(20*2)

P(Red AND Tails) = 12/40

P(Red AND Tails) = 3/10
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------